# Why CAT(0) cube complexes should be replaced with median graphs

### v2

@@ Abstract
Interestingly, the one-skeletons of CAT(0) cube complexes coincide with median graphs, a well-known family of graphs, formally introduced in the 1960s [Ava61, Neb71] but with even older roots (see for instance [BC08, Mul11] and references therein).
@@
Very few articles actually use CAT(0) geometry in an essential way to prove results about CAT(0) cube complexes. Instead, they mainly use the graph-metric and the combinatorics of hyperplanes. Nevertheless, before replacing “CAT(0) cube complexes” with “median graphs”, at the very least we have to verify that the central results about the structure of CAT(0) cube complexes and about group actions used in geometric group theory can be easily and naturally proved by using the point of view of median graphs. This approach is the leitmotiv of the project [Gen].
Personally, I am not aware of a single theorem about CAT(0) cube complexes that cannot be naturally proved thanks to median graphs.
@@
So far, our goal was to promote the use of “median graphs” instead of “CAT(0) cube complexes”. However, another alternative, not mentioned so far, could be “median cube complexes”, in order to keep the higher-dimensional cellular structure. Here, a median cube complex would refer to a cube complex such that the length metric extending the $\ell^{1}$-metrics on each cube (whose sides, by convention, have length one) is median.

### v3

@@ Abstract
Interestingly, the one-skeletons of CAT(0) cube complexes coincide with a well-known family of graphs: median graphs. Formally introduced in the 1960s-70s [Ava61, Neb71, MS79], but with even older roots motivated by lattice theory, median graphs have been extensively investigated in metric graph theory. (See for instance [BC08, Mul11] and references therein.)
@@
Very few articles actually use CAT(0) geometry in an essential way to prove results about CAT(0) cube complexes. Instead, they mainly use the graph-metric and the combinatorics of hyperplanes. Nevertheless, before replacing “CAT(0) cube complexes” with “median graphs”, at the very least we have to verify that the central results about the structure of CAT(0) cube complexes and about group actions used in geometric group theory can be easily and naturally proved by using the point of view of median graphs. This is done, for instance, in [Gen].
In order to justify the replacement of CAT(0) cube complexes with median graphs, we do not have to prove that CAT(0) geometry will never be relevant. Actually, there may (and probably do) exist a few specific cases where CAT(0) geometry on cube complexes is superior from some perspective [^1] Personally, I am not aware of a single theorem about CAT(0) cube complexes that cannot be naturally proved thanks to median graphs.. But it suffices to show, as tried in this note and in [Gen], that median geometry is more natural, efficient, and satisfying in most cases.
@@
§6 Historical background.
One could argue that it would be reasonable to keep the terminology “CAT(0) cube complexes” in order to keep track of the historical roots of the objects. But this would be a point of view restricted to geometric group theory, ignoring the perspective offered by metric graph theory. Median graphs have an even older history in metric graph theory. Highlighting this history would be helpful in order to promote interactions between geometric group theory and metric graph theory. Despite the fact that many concepts related to CAT(0) cube complexes in geometric group theory can be found in metric graph theory several years earlier [^6] A notable example is provided by Buneman graphs of cut systems [Bar89], which turn out to coincide with cubulations of wallspaces [Nic04, CN05]., very few geometric group theorists have some background in metric graph theory.
As an illustration of the benefit that can follow from the perspective offered by metric graph theory, let us mention the Djoković-Winkler relation, which sheds an interesting light on the notion of hyperplanes. Given a graph $X$, say that two edges $\{a,b\}$ and $\{x,y\}$ are in $\theta$-relation if
$d(a,x)+d(b,y)\neq d(a,y)+d(b,x).$
Alternatively, the edge $\{a,b\}$ yields a partition of $X$ into three vertices: the vertices closer to $a$ than to $b$, the vertices closer to $b$ than to $a$, and the vertices at equal distance to $a$ and $b$. The edge $\{x,y\}$ is then in $\theta$-relation with $\{a,b\}$ whenever $x$ and $y$ belong to two distinct pieces of the partition.
The Djoković-Winkler relation is reflexive and symmetric, but it may not be transitive. So we let $\theta^{\ast}$ denote its transitive closure.
The $\theta$-relation was introduced in [Djo73], and latter generalised in [Win84] for non-bipartite graphs, in order to characterise partial cubes, i.e. isometrically embedded subgraphs in hypercubes. Then, [GW85] proved that the relation plays a fundamental role in representations of graphs as subgraphs of Cartesian products. More precisely, given a graph $X$, there is a canonical and optimal isometric embedding
$X\hookrightarrow\prod\limits_{e\in E(X)/\theta^{\ast}}X//e^{c}$
where each $X//e^{c}$ denotes the graph obtained from $X$ by collapsing all the edges not in the class $e$.
When applied to median graphs, $\theta^{\ast}$ can be described as the reflexive-transitive closure of the relation that identifies two edges whenever they are opposite in some $4$-cycle. In other words, we recover the standard notion of hyperplanes. And the embedding above coincides with the well-known embedding of a median graph inside a hypercube (whose will be, in addition, a retract). Hence a much deeper perspective on the notion of hyperplanes in CAT(0) cube complexes, which echoes some constructions of cubulations that implicitly use the Djoković-Winkler relation (e.g. Coxeter groups [NR03]).
@@
So far, our goal was to promote the use of “median graphs” instead of “CAT(0) cube complexes”. However, other alternatives, not mentioned so far, exist. For instance, “median cube complexes”, in order to keep the higher-dimensional cellular structure. Here, a median cube complex would refer to a cube complex such that the length metric extending the $\ell^{1}$-metrics on each cube (whose sides, by convention, have length one) is median.